The probability of a man hitting a target is 1/4. He fires 6 times. Find the probability that he hits the target at least once.


1
Expert's answer
2021-03-02T04:47:24-0500

We have that

P(hitting a target) = 1/4 = 0.25

n = 6

Need to find "P(X\ge1) = 1-P(X=0)"

This follows binomial distribution.

The binomial probability is calculated by the formula:


"P(X=m)=C(n,m)\cdot p^m \cdot (1-p)^{n-m}"

where m = 0

"P(X=0)=C(6,0)\cdot 0.25^0 \cdot (1-0.25)^{6-0}=\frac{6!}{0!6!}\cdot0.25^0\cdot0.75^{6}=0.178"

"P(X\ge1) = 1-P(X=0)=1-0.178=0.822"


Answer: 0.822